OpenStudy (anonymous):

What is the slope and y-intercept of 7x–1=3y+8

OpenStudy (haleyelizabeth2017):

you need to turn it into the form of y=mx+b, can you do that?

OpenStudy (anonymous):

go to mathway.com

OpenStudy (haleyelizabeth2017):

@gummybear1021 it will be easier just to go through it on here so @Mathwizard24-7 will know what to do for the next time :)

OpenStudy (texaschic101):

7x - 1 = 3y + 8 put this in y = mx + b form 7x - 1 - 3y = 8 -3y = -7x + 1 + 8 -3y = -7x + 9 y = 7/3x - 3

OpenStudy (anonymous):

@haleyelizabeth2017 yes but if ge or she needs a quick answer they will get it

OpenStudy (haleyelizabeth2017):

That is true.

OpenStudy (texaschic101):

slope is in the m position and the y intercept is in the b position

OpenStudy (texaschic101):

so the slope is 7/3 and the y int is -3